OpenStudy (anonymous):

ASAP a basketball player standing under the hoop shoots the ball straight up with an initial velocity of v=3.5 from 2.5m above the ground. What is the max height, above the launch point the basketball will achieve? a formula provided would be nice.

OpenStudy (anonymous):

vf^2=vi^2 +2gh <----- is that the right equation for this problem?

OpenStudy (anonymous):

s=ut + 1/2at^2 and here a=g with a minus sign as ball is retarding